Gram equivalent volume of O2 at STP is? Plz provide proper solution.

Grade:12th pass

2 Answers

Arun
4 years ago
it goes like this, 
oxygen atomic mass=16 grams 
valency=2 
oxygen equivalent mass=8(16/2) 
now,at NTP 32 grams of oxygen=22.4 litres 
8 grams=22.4/32 multiplied by 8 
=5.6
RIDA HASNAIN
one year ago
 O2 + 4e- ------- 2O2-
1 mole of o2 at STP occupied 22.4L of space 
4g - equivalent of o2 at STP occupied 22.4 L of space
1g equivalent of o2 at STP =22.4 /4 
                                              =5.6